OpenStudy (anonymous):

What is the molarity of a solution that 85.0 grams of Na2SO4 in 325 ml of solution

OpenStudy (anonymous):

I literally just answered a question like this. Molarity = \[\frac{ Moles of Solute }{ Liters of Solution}\] There are 142 grams in a mole of Na2SO4. \[\frac{ 85.0g }{ 142 g/mol } = 0.59859154929 moles\] 325 milliliters of solution is equal to .325 Liters. Putting this back into the equation we get Molarity =\[\frac{ 0.59859154929 }{ .325 } = 1.84182015166\] But we are limited to 3 significant figures, so the answer would be 1.84M